# Low-resource Machine Translation: what for? who for? An observational study on a dedicated Tetun language translation service

### v1

@@ Abstract
The impact of machine translation (MT) on low-resource languages remains poorly understood. In particular, observational studies of actual usage patterns are scarce. Such studies could provide valuable insights into user needs and behaviours, complementing survey-based methods. Here we present an observational analysis of real-world MT usage for Tetun, the lingua franca of Timor-Leste, using server logs from a widely-used MT service with over 70,0007000070,00070 , 000 monthly active users. Our analysis of 100,000100000100,000100 , 000 translation requests reveals patterns that challenge assumptions based on existing corpora. We find that users, many of them students on mobile devices, typically translate short texts into Tetun across diverse domains including science, healthcare, and daily life. This contrasts sharply with available Tetun corpora, which are dominated by news articles covering government and social issues. Our results suggest that MT systems for languages like Tetun should prioritise translating into the low-resource language, handling brief inputs effectively, and covering a wide range of domains relevant to educational contexts. More broadly, this study demonstrates how observational analysis can inform low-resource language technology development, by grounding research in practical community needs.
@@
Raphael Merx1, Hanna Suominen2, Adérito José Guterres Correia3, Trevor Cohn1, Ekaterina Vylomova1 1School of Computing and Information Systems, The University of Melbourne 2The Australian National University 3Instituto Nacional de Linguística, Dili rmerx@student.unimelb.edu.au
@@ 1 Introduction
While machine translation (MT) is often considered a single area for natural language processing (NLP), it covers a large range of applications, end-user requirements, and language directions. MT can be leveraged for everyday communication, learning, or translation of official communication Dew et al. (2018); Vieira et al. (2021); Herrera-Espejel and Rach (2023); Paterson (2023). It can be used by students in the classroom Deng and Yu (2022), workers in an office environment Brynjolfsson et al. (2019), or by foreign travellers in public spaces Carvalho et al. (2023). Its reliability varies significantly depending on the size and domain coverage of available corpora Dabre et al. (2020); Khiu et al. (2024), and similarity between source and target languages Xu et al. (2020).
Figure 1: Real MT usage collides with assumptions made from available corpora
This disparity of potential applications and accuracy across domains or language pairs is particularly salient when translating to or from low-resource languages Haddow et al. (2022). In low-resource scenarios, the combination of high variability in MT accuracy across domains Haddow et al. (2022), and high variability of what native speakers value in NLP applications Lent et al. (2022), has prompted NLP researchers to better anchor their work in local community perception, including through surveys, focus groups, and product design Mager et al. (2023); Blaschke et al. (2024); Le Ferrand et al. (2022). These approaches, however, could be complemented by observational studies, that rely on end-user behaviours to infer needs or lack thereof for MT across domains or demographics.
In this paper on low-resource MT, we work with the Tetun language, which is the most spoken language in Timor-Leste General Directorate of Statistics Timor-Leste (2015). We rely on a sample of server logs from a volunteer-run MT service for the Tetun language, available in both website and app format. The service has over 70,0007000070,00070 , 000 monthly active users, and translates over one million documents per month. From a sample of 100,000100000100,000100 , 000 end-user translation text inputs, and their associated machine-translated output, we aim to respond to the following research questions:
  • Who are the main users of this service, what technology do they rely on, and what is the driver of their MT usage?
  • What characterises MT inputs, in terms of domain, length, direction of translation, and language pairs?
@@ 2 Related Work
Tetun and the NLP community. While Tetun is a low-resource language de Jesus (2024), as of the 2015 census, it was the largest mother tongue in Timor-Leste with 361,027 speakers, and an additional 571,406 speakers of Tetun as a second language General Directorate of Statistics Timor-Leste (2015). It is the lingua franca in this country, spoken in most public spaces, and in the media Greksakova (2018). To our knowledge, the majority of Tetun NLP research to date was published by Gabriel de Jesus during his PhD candidature at the University of Porto. This included the creation of a language identification model, language tokenizer, and data pipeline for gathering Tetun monolingual data de Jesus and Nunes (2024a), the publication of that corpus de Jesus and Nunes (2024c), the study of retrieval strategies for Tetun text de Jesus (2024), and the study of different large language models (LLMs) relevance judgement capabilities in Tetun de Jesus and Nunes (2024b).
Community-driven approaches to NLP for low-resource languages. Recent years have witnessed a growing focus on developing NLP technologies for low-resource and endangered languages. However, researchers have identified that academic motivations can be at odds with the actual needs of language communities Le Ferrand et al. (2022); Liu et al. (2022), which made them advocate for a paradigm shift towards more community-centered approaches in NLP research Lent et al. (2022); Bird and Yibarbuk (2024). Studies emphasise the importance of close collaboration with language communities, understanding their specific needs, and involving them in the design and development process of language technologies Mager et al. (2023); Bird and Yibarbuk (2024). This approach acknowledges the heterogeneity among low-resource languages and their communities, recognising that methodologies effective for one may not be universally applicable Lent et al. (2022); Dolinska et al. (2024).
Linguistic analysis through large-scale data gathering. To our knowledge, this paper is the first large-scale analysis of real-world usage patterns in MT. It takes a similar approach to prior work leveraging large-scale secondary data in linguistics, including those predicting language learning difficulty Schepens et al. (2020), understanding what language characteristics lead to a language being confused for another Skirgård et al. (2017), and games that aim to incentivise the development of new language resources Fiumara et al. (2022).
@@ 3 Methodology
We work with data from a specialised Tetun MT service, which allows automated translation between Tetun Dili (tdt) and any of English (eng), Indonesian (ind), or Portuguese (por). The service also includes bilingual word dictionaries and a spell checker for the Tetun language, but we focus on the MT part. We get access to the following data sources:
@@
    A sample of 100,000100000100,000100 , 000 server logs from the machine translation service, all collected between March and August 2024. For each log, we get data on timestamp, translation input, MT output, source language code, requested target language code, and device OS.
@@
    Summary monthly active users’ data by country and device sourced from Google Analytics (for the website), Google Play Store (for the Android app), and Apple App Store (for the iOS app).
While data on the user base can be directly extracted from (2), domain analysis requires a choice of methodology that we describe in the following subsection.
@@ 3.1 Domain classification
Components of domain
Recognising that "domain" is often used interchangeably in NLP research to refer to topic (e.g. "the health domain") or source (e.g. "the news domain"), we follow the methodology in Saunders (2022), which defines domain as a combination of topic (one or more subject matters covered by the text), provenance (where the text comes from), and genre (language syntax and style).
@@ 3.1.1 Topic classification
We start by identifying topics using Latent Dirichlet Allocation (LDA, Blei et al., 2003) on documents that are more than 5 words long, as very short documents do not offer enough context to identify their topic. We work with the Tetun side (either source Tetun text, or target machine-translated Tetun text), as this language is always present in any of the language pairs. The text is first lower cased and punctuation is removed. We run LDA enhanced with bigrams for improved context. After running hyperparameter search on number of topics, passes, and alpha, we find a maximum topic coherence of 0.065 using 15 topics, 10 passes, and an asymetric alpha. While this informs our choice of topic to categorise from, we qualitatively find that LDA is noisy for this task. For example, the health topic also includes words related to project management, such as “objetivu” or “servisu”. It also performs poorly on the categorisation of shorter documents, which are more likely to contain words that are not part of any identified topics.
Labels & test set preparation
@@
Through a manual review of words for each of the LDA topic, we land on the following list of topics to classify from: Technology, Daily Life & Personal Experiences, Healthcare & Medicine, Law & Regulations, Business & Work & Employment, Government & Socio-economic Issues, Education, Science & Research, Religion & Spirituality. We map each of the 15 LDA topics to one of those topics, and we manually annotate a sample of 100 logs against this list of topics for evaluation, where each log can be classified against multiple topics.
Classification using pre-trained models
@@
We evaluate other classification methods that rely on pre-trained models: zero-shot transformer-based classifiers and LLMs. Given the lack of pre-trained models that cover Tetun, we rely on the high-resource side of the input-output pair, which can be in English, Indonesian or Portuguese. For zero-shot classifiers, we use the HuggingFace zero shot classification pipeline [^1] https://huggingface.co/tasks/zero-shot-classification with the list of topics passed as candidate labels, in a multi-label mode. We evaluate bart-large-mnli [^2] https://huggingface.co/facebook/bart-large-mnli and deberta-v3-large-zeroshot-v2.0333https://huggingface.co/MoritzLaurer/deberta-v3-large-zeroshot-v2.0 Laurer and Lochner (2024). For LLMs, we evaluate LLama 3.1 8b Dubey et al. (2024), with a prompt that lists the domains to classify from, and that includes examples for in-context learning, shown in Appendix A.
Classification evaluation
@@
We compare micro and weighted average scores on our test set for the different classification approaches in Table 1. We find that Llama 3.1 8B has the highest F1 score on both metrics for this classification task, followed by (in order) deberta-v3-large-zeroshot-v2.0, bart-large-mnli, and last LDA. We therefore pick Llama 3.1 8B for domain classification, with results reported in Section 4.3.
@@
bart-mnli4442 $0.400.40$ $0.400.40$
deberta-nli5553 $0.660.66$ $0.680.68$
@@
Table 1: Comparison of micro and weighted average F1 score for different domain classification approaches. LDA is run on the Tetun side, while other approaches are run on the high-resource side (English/Portuguese/Indonesian).
@@ 3.1.2 Provenance classification
After a manual review of the text samples used in topic classification, we propose the following categories for provenance: Education & research material, News article & press release, Organisational & formal documents, Conversation & correspondence, Literary fiction, Religious text, Other. We use Llama 3.1 8B here as well, with a prompt that lists the sources to classify from, and that includes examples for in-context learning. This approach results in a weighted F1 score of 0.73 on our test set, which we deem satisfactory for classifying provenance, given the natural ambiguity of this task, especially for shorter documents. We report our results in Section 4.3.2.
@@ 3.1.3 Genre
After recommendation from the Timor-Leste Instituto Nacional de Linguística (INL, National Institute of Linguistics), under genre we evaluate what fraction of Tetun MT inputs follow official spelling, as defined by the country’s Matadalan Ortográfiku ba Tetun Nasionál (Orthographic Guide for National Tetun, 2002).
For calculating the percentage of words that follow INL spelling in each document, we apply the following preprocessing: (1) tokenize Tetun document using an open source Tetun tokenizer de Jesus and Nunes (2024a), (2) ignore words that start with a capital letter as they are often proper nouns and (3) ignore tokens that start with a number or punctuation character. We report our results in Section 4.3.3. We further ignore inputs where less than 20% of words are in the official spelling guide, as we empirically find that these documents are almost never in Tetun.
@@ 3.1.4 Comparison with available monolingual corpora
For each of the identified domain components (topic, provenance, genre), we provide a comparison between MT inputs and available Tetun corpora. For the latter, we select the Tetun "clean" split of the MADLAD-400 dataset Kudugunta et al. (2023), as it is the largest Tetun corpus available to date with 40,3674036740,36740 , 367 documents. For classifying topic and provenance, given that we work with models that do not support Tetun, we first translate the MADLAD Tetun corpus to English using the MADLAD-3B model. We release our translation of this corpus for future researchers to use, which includes 884k Tetun sentences and their machine-translated text.
For all domain analysis, we weigh inputs by their number of words. This is to ensure that very short MT inputs do not get as much weight as longer, more informative inputs, or that users that tend to translate several paragraphs at a time are not under-represented compared to users who translate shorter inputs.
@@ 4.1 User base
Device. Most users access the service through their mobile device, either through the Android application (71,5007150071,50071 , 500 monthly active users), the iOS application (2,50025002,5002 , 500 monthly active users), or the website in a mobile browser (13,0001300013,00013 , 000 monthly active users). Desktop users of the website represented 8,30083008,3008 , 300 monthly active users, with an unknown number of these users being duplicate of mobile users. The distribution of users is illustrated in Figure 2. We discuss the implications of this high representation of mobile devices, particularly Android devices, on NLP research, in Section 5.
MobileDesktop00 k20202020 k40404040 k60606060 k80808080 kGoogle Play StoreApple App StoreWebsite
@@
Country. Most users are in Timor-Leste: 67,000 of 71,500 (93.4%) Android monthly active users have their device registered in Timor-Leste, and 17,000 of 22,000 (77%) website monthly active users are in Timor-Leste. These numbers can be noisy: for example, Timorese users who live abroad might have a mobile device registered outside of Timor-Leste, and foreigners in Timor-Leste using the website cannot be differentiated from local users. We think, however, that the results support the hypothesis that the large majority of users are Timorese with English as a second language (ESL), as opposed to foreigners learning Tetun.
Demographics. Server usage patterns support the hypothesis that most users are students who use the service for educational purposes. A large proportion of requests happen in the early morning or in the evening (see Figure 3), at times when office workers are less likely to use the service, and when students are more likely to do their homework. Authors of the MT service also communicated to us that service usage goes down during school holidays, and back up on the evening before school resumes, another pattern that supports this hypothesis.
004444888812121212161616162020202000 k0.50.50.50.5 k1111 k1.51.51.51.5 k2222 kHour of day
Figure 3: Histogram showing number of requests by time of day, in Asia/Dili (UTC+9). We observe a spike in the evenings, when students are more likely to prepare their homework.

### v4

@@ Abstract
Low-resource machine translation (MT) presents a diversity of community needs and application challenges that remain poorly understood. To complement surveys and focus groups, which tend to rely on small samples of respondents, we propose an observational study on actual usage patterns of tetun.org, a specialized MT service for the Tetun language, which is the lingua franca in Timor-Leste. Our analysis of 100,000100000100,000100 , 000 translation requests reveals patterns that challenge assumptions based on existing corpora. We find that users, many of them students on mobile devices, typically translate text from a high-resource language into Tetun across diverse domains including science, healthcare, and daily life. This contrasts sharply with available Tetun corpora, which are dominated by news articles covering government and social issues. Our results suggest that MT systems for institutionalized minority languages like Tetun should prioritize accuracy on domains relevant to educational contexts, in the high-resource to low-resource direction. More broadly, this study demonstrates how observational analysis can inform low-resource language technology development, by grounding research in practical community needs.
@@
Raphael Merx1, Adérito José Guterres Correia2, Hanna Suominen3, Ekaterina Vylomova1 1School of Computing and Information Systems, The University of Melbourne 2Instituto Nacional de Linguística, Dili 3The Australian National University rmerx@student.unimelb.edu.au
@@ 1 Introduction
While machine translation (MT) is often considered a single area for natural language processing (NLP), it covers a large range of applications, end-user requirements, and language directions. MT can be leveraged for everyday communication, specialized sectors (e.g., health or education), learning, or translation of official communication Dew et al. (2018); Vieira et al. (2021); Herrera-Espejel and Rach (2023); Paterson (2023); Merx et al. (2024). It can be used by students in the classroom Deng and Yu (2022), workers in an office environment Brynjolfsson et al. (2019), or by foreign travelers in public spaces Carvalho et al. (2023). Its reliability varies significantly depending on the size and domain coverage of available corpora Dabre et al. (2020); Khiu et al. (2024), and depending on similarity between source and target languages Xu et al. (2020).
Figure 1: Overview of our approach and research questions
This disparity of potential applications and accuracy across domains or language pairs is particularly salient when translating to or from low-resource languages Haddow et al. (2022). In low-resource scenarios, the combination of high variability both in MT accuracy across domains Haddow et al. (2022) and of what native speakers value in NLP applications Lent et al. (2022), has prompted NLP researchers to better anchor their work in local community perception, including through surveys, focus groups, and product design Le Ferrand et al. (2022); Mager et al. (2023); Blaschke et al. (2024). These approaches, however, could be complemented by observational studies that rely on inferring needs or lack thereof for MT across domains or demographics from end-user behaviors.
In this paper on low-resource MT, we work with the Tetun language, [^1] Also called Tetun Dili, or Tetum. which is the most spoken language in Timor-Leste General Directorate of Statistics Timor-Leste (2015). We rely on a sample of server logs from tetun.org, a volunteer-run MT service for the Tetun language, available in both website and app222play.google.com/store/apps/details?id=org.tetun.tetunorg format. Until Google Translate added Tetun support in June 2024, this service was the only widely available MT tool for Tetun. Tetun.org has over 70,0007000070,00070 , 000 monthly active users, and translates over one million documents per month. From a sample of 100,000100000100,000100 , 000 end-user translation text inputs, and their associated machine-translated output, collected between March and August 2024, we aim to respond to the following research questions (Figure 1):
  • Who are the main users of tetun.org, what technology do they rely on, and what is the driver of their MT usage?
  • What characterizes MT inputs, in terms of domain, length, direction of translation, and language pairs?
@@
Tetun: low-resource, but institutional.
Following Nigatu et al. (2024), we define why Tetun is low-resource in our view: (1) Tetun has very limited digital resources, with less than one million monolingual sentences available in open corpora de Jesus and Nunes (2024c) (2) Tetun only has around one million native speakers General Directorate of Statistics Timor-Leste (2015) and (3) Tetun has very little NLP research coverage, apart from work by Gabriel de Jesus during his PhD candidature at the University of Porto de Jesus and Nunes (2024a); de Jesus (2024); de Jesus and Nunes (2024b). However, Tetun is an institutional language, with both official status in Timor-Leste, and active usage in education and the media Greksakova (2018). Our findings, while focused on Tetun, may extend to other institutionalized low-resource languages, but are less likely to apply to non-institutionalized or endangered ones.
@@ 2 Related Work
Community-driven approaches to NLP for low-resource languages.
Recent years have witnessed a growing focus on developing NLP technologies for low-resource and endangered languages. Researchers have identified that academic motivations can be at odds with the actual needs of language communities Le Ferrand et al. (2022); Liu et al. (2022), which made them advocate for a paradigm shift towards more community-centered approaches in NLP research Lent et al. (2022); Bird and Yibarbuk (2024). Studies emphasize the importance of close collaboration with language communities, understanding their specific needs, and involving them in the design and development process of language technologies Mager et al. (2023); Bird and Yibarbuk (2024). This approach acknowledges the heterogeneity among low-resource languages and their communities, recognizing that methodologies effective for one may not be universally applicable Lent et al. (2022); Dolinska et al. (2024).
Linguistic analysis through large-scale data gathering.
To our knowledge, this paper is the first large-scale analysis of real-world usage patterns in MT. It takes a similar approach to prior work leveraging large-scale secondary data in linguistics, including those predicting language learning difficulty Schepens et al. (2020), understanding what language characteristics lead to a language being confused for another Skirgård et al. (2017), and games that aim to incentivise the development of new language resources Fiumara et al. (2022).
Domain imbalance and low-resource MT
Low-resource language corpora are often concentrated in specific domains, particularly religious texts Marashian et al. (2025), and lack the broad domain coverage typically found in their high-resource counterparts Haddow et al. (2022). As a consequence, NLP tools trained on these available corpora tend to lack domain robustness Ranathunga et al. (2023). This issue is particularly significant for MT systems, as demonstrated by Khiu et al. (2024), who find that domain similarity between training and test data is the largest driver of low-resource MT performance, more so than training data size or language similarity between source and target languages. Building on these findings, our analysis pays particular attention to differences of domain distribution between available corpora and actual translation requests.
@@ 3 Methodology
We work with data from tetun.org, a specialized service for automated translation between Tetun Dili (tdt) and any of English (eng), Indonesian (ind), or Portuguese (por). The service also includes bilingual word dictionaries and a spell checker for the Tetun language, but we focus on the MT part.
The MT service uses a transformer model, and inference is run on the server side. The system was initially developed by fine-tuning a M2M100 multilingual model Fan et al. (2020) on a scraped parallel Tetun-English corpus. To optimize inference speed, this model was then distilled into a smaller transformer-base architecture. For Portuguese and Indonesian translation, the system uses English as a pivot language. The model’s performance was evaluated using BLEU scores on a curated test dataset assembled by tetun.org volunteers.
We get access to the following data sources:
@@
    A sample of 100,000100000100,000100 , 000 server logs from the service, all collected between March and August 2024. Each log contains data on timestamp, text input, MT output, source language code, requested target language code, and device operating system (OS).
@@
    Summary monthly active users’ data by country and device sourced from Google Analytics (for the website), Google Play Store,333play.google.com/store/apps/details?id=org.tetun.tetunorg and Apple App Store.444apps.apple.com/us/app/tetun-org/id1515892208
While data on the user base can be directly extracted from (2), domain analysis requires a choice of methodology that we describe next.
@@ 3.1 Domain classification
Components of domain.
Recognizing that “domain” is often used interchangeably in NLP research to refer to topic (e.g., “the health domain”) or source (e.g., “the news domain”), we follow the methodology in Saunders (2022), which defines domain as a combination of topic (one or more subject matters covered by the text), provenance (where the text comes from), and genre (language syntax and style).
@@ 3.1.1 Topic classification
We start by identifying topics using Latent Dirichlet Allocation (LDA, Blei et al., 2003) on documents that are more than 5 words long, as very short documents do not offer enough context to identify their topic. We work with the Tetun side (either source Tetun text, or target machine-translated Tetun text), as this language is always present in any of the language pairs. The text is first lower-cased and punctuation is removed. We run LDA enhanced with bigrams that appear at least 5 times in the corpus for improved context. After running hyperparameter search on number of topics, passes, and alpha, we find a maximum topic coherence of 0.065 using 15 topics, 10 passes, and an asymmetric alpha. While this informs our choice of topic to categorize from, we qualitatively find that LDA is noisy for this task. For example, the health topic also includes words related to project management, such as “objetivu” or “servisu”. It also performs poorly on the categorization of shorter documents, which are likely to contain words that are not part of any identified topics. However, we leverage these identified topics to formulate a list of topics to classify from.
Labels & test set preparation.
@@
Through a manual review of each LDA topic, we land on the following list of topics to classify from: Technology, Daily life & personal experiences, Healthcare & medicine, Law & regulations, Business & work & employment, Government & socio-economic issues, Education, Science & research, Religion & spirituality. We map each of the 15 LDA topics to one of those topics, and our first author manually annotates a sample of 100 logs against this list of topics for evaluation, where each log can be classified against multiple topics.
Classification using pre-trained models.
@@
We evaluate other classification methods that use pre-trained models, either zero-shot transformer-based classifiers or LLMs. Given the lack of pre-trained models that cover Tetun, we rely on the high-resource side of the input-output pair, which can be in English, Indonesian or Portuguese. For zero-shot classifiers, we use the HuggingFace zero shot classification pipeline555huggingface.co/tasks/zero-shot-classification with the list of topics passed as candidate labels, in a multi-label mode. We evaluate bart-large-mnli666huggingface.co/facebook/bart-large-mnli and deberta-v3-large-zeroshot-v2.0777huggingface.co/MoritzLaurer/deberta-v3-zeroshot-v2.0 Laurer et al. (2023). For LLMs, we assess LLama 3.1 8B Dubey et al. (2024), with a prompt that lists the domains to classify from, and that includes examples for in-context learning, shown in Appendix A.
Classification evaluation.
@@
We compare micro and weighted average scores on our test set for the different classification approaches in Table 1. We find that Llama 3.1 8B has the highest weighted averaged F1 score on both metrics for this classification task, followed by (in order) deberta-v3-large-zeroshot-v2.0, bart-large-mnli, and last LDA.
We therefore pick Llama 3.1 8B for topic classification, and run topic inference on the whole 100,000100000100,000100 , 000 server logs, using a single Nvidia A100 GPU, which takes approximately 30 hours.
@@
bart-mnli $0.400.40$ $0.400.40$
deberta-nli $0.660.66$ $0.680.68$
@@
Table 1: Comparison of micro and weighted averaged F1 score for different domain classification approaches. LDA is run on the Tetun side, while other approaches are run on the high-resource side (English/Portuguese/Indonesian).
@@ 3.1.2 Provenance classification
After a manual review of the text samples used in topic classification, we propose the following categories for provenance: Education & research material, News article & press release, Organizational & formal documents, Conversation & correspondence, Literary fiction, Religious text, Other. We use Llama 3.1 8B here as well, with a prompt that lists the sources to classify from, and that includes examples for in-context learning. On the same sample of 100 logs, which we manually annotate for provenance, we get a weighted F1 score of 0.73, which we deem satisfactory for classifying provenance, given the natural ambiguity of this task, especially for shorter documents.
@@ 3.1.3 Genre
After recommendation from the Timor-Leste Instituto Nacional de Linguística (INL, National Institute of Linguistics), under genre, we evaluate what fraction of Tetun MT inputs follow official spelling, as defined by the country’s Matadalan Ortográfiku ba Tetun Nasionál (Orthographic Guide for National Tetun, 2002).
For calculating the percentage of words that follow INL spelling in each document, we apply the following preprocessing: (1) tokenize Tetun document using an open source Tetun tokenizer de Jesus and Nunes (2024a), (2) ignore words that start with a capital letter as they are often proper nouns and (3) ignore tokens that start with a number or punctuation character. We further ignore inputs where less than 20% of words were in the official spelling guide, as we empirically find that these documents are almost never in Tetun.
@@ 3.1.4 Comparison with available monolingual corpora
For each of the identified domain components (topic, provenance, genre), we provide a comparison between MT inputs and available Tetun corpora. For the latter, we select the Tetun “clean” split of the MADLAD-400 dataset Kudugunta et al. (2023), as it is the largest Tetun corpus available to date with 40,3674036740,36740 , 367 documents. For classifying topic and provenance, given that we work with models that do not support Tetun, we first translate the MADLAD Tetun corpus to English using the MADLAD-3B model. Upon acceptance, we will release our translation of this corpus of 884,000884000884,000884 , 000 Tetun sentences and their machine-translated text for future researchers to use.
For all domain analysis, we weight inputs by their number of words. This is to ensure that very short MT inputs do not get as much weight as longer, more informative inputs, or that users who tend to translate several paragraphs at a time are not under-represented compared to users who translate shorter inputs.
@@ 4.1 User base
Country. Most users are in Timor-Leste: 67,000 of 71,500 (93.4%) Android monthly active users have their device registered in Timor-Leste, and 17,000 of 22,000 (77%) website monthly active users are in Timor-Leste. These results support the hypothesis that the large majority of users are Timorese with English as a second language (ESL), as opposed to foreigners learning Tetun. While usage numbers can be noisy (for example, Timorese users who live abroad might have a mobile device registered outside of Timor-Leste, and foreigners in Timor-Leste using the website cannot be differentiated from local users), given the larger number of Timorese emigrating than people immigrating to Timor (INETL, INETL), they are more likely to under-estimate than to over-estimate the proportion of Timorese users.
Device. Most users access tetun.org through their mobile device, either through the Android application (71,5007150071,50071 , 500 monthly active users), the iOS application (2,50025002,5002 , 500 monthly active users), or the website in a mobile browser (13,0001300013,00013 , 000 monthly active users). Desktop users of the website represent 8,30083008,3008 , 300 monthly active users, with an unknown number of these users being duplicate of mobile users. The distribution of users is illustrated in Figure 2. We discuss the implications of this high representation of mobile devices, particularly Android devices, on NLP research, in Section 5.
MobileDesktop00 k20202020 k40404040 k60606060 k80808080 k13.613.613.613.6% 8.78.78.78.7% 2.62.62.62.6% 75.075.075.075.0% Google Play StoreApple App StoreWebsite
@@
Demographics. Server usage patterns seem to indicate that most users are students who use tetun.org for educational purposes. This is supported by (1) the large proportion of requests that occur during evenings before school days (including on Sundays, see Figure 3), at times when office workers are less likely to use the service, and when students are more likely to do their homework; (2) the decrease in service usage during school holidays, with usage going back up the evening before school resumes; (3) domain analysis results presented in Section 4.3.
Figure 3: Number of requests by day / hour, in Asia/Dili (UTC+9). We observe a spike in the evenings before school days (Saturday is a school day in Timor-Leste), when students are more likely to prepare their homework.